OpenStudy (anonymous):

You find the computer of your dreams. When you go to pay for it, the total price is $958. This was after 1/5 of the total price was added for insurance and taxes. A. Using c for computer, write an algebraic equation to represent the problem. (5 points) B. Using algebraic transformations, solve the equation you wrote for Part A. Show all steps. (10 points) C. How much is the computer before taxes and insurance? Remember to label your answer

OpenStudy (anonymous):

@texaschic101

OpenStudy (anonymous):

958 = (1.2) c

OpenStudy (anonymous):

@douglaswinslowcooper can u label the answers a b and c

OpenStudy (anonymous):

sorry. must go that was for first question, a

OpenStudy (texaschic101):

c + 1/5(958) = 958 (A) c + 958/5 = 958 --- (B) c + 191.60 = 958 c = 958 - 191.60 c = 766.40 ----(C)